OpenStudy (anonymous):

Find the value of angle c. Explain/justify your steps a= 2x + 10 b= 3x +20

OpenStudy (tkhunny):

Is it a triangle? The entire problem statement would help. Showing your work would help even more.

OpenStudy (anonymous):

no it is not a triangle i uploaded the attachment above ..

OpenStudy (anonymous):

a+b = 180 a = c

OpenStudy (anonymous):

how did you get that answer ?

OpenStudy (anonymous):

@swagmaster47

OpenStudy (tkhunny):

Some folks will not open a .DOC file.

OpenStudy (anonymous):

can you help me please ? @tkhunny

OpenStudy (tkhunny):

Not without the rest of the problem. I happen to be one of those who will not open a .DOC file.

OpenStudy (anonymous):

why ..

OpenStudy (anonymous):

Are you serious?

OpenStudy (anonymous):

Can someone please help me, I will give you a medal and become a fan , someone please.

OpenStudy (anonymous):

@adorablefourfootnine some people don't want to because they think it's going to be inappropriate

OpenStudy (anonymous):

But I will c if I can help you. :)

OpenStudy (anonymous):

I don't know what you mean by that ? and I'd greatly appreciate it

OpenStudy (anonymous):

Angle A and angle C are the same. Did u notice that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how

OpenStudy (anonymous):

If you look back at the word document, you'll see that angle a and angle c have the same angles. Therefore, they're the same.

OpenStudy (anonymous):

yeah I did already, but what does that have to do with the equations i just put up?

OpenStudy (anonymous):

sup

OpenStudy (anonymous):

If: a= 2x + 10 b= 3x +20 Then: a+b=180 (sum of angles on straight line add up to 180) 2x+10+3x+20=180 5x+30=180 5x=180-30 5x=150 (5x/5)=(150/5) x=30  a=c (vertically opposite angles) c=30

OpenStudy (anonymous):

Thank you! @Rudi4Life